A class of twenty four pupils consists of 11 girls and 13 boys. To form the class committee, four of the pupils are chosen at random as "Chairperson", "Vice Chairperson", "Treasurer" and "Secretary".

  1. (i)

    Find the probability that the committee will consist of at least one girl and at least one boy.

  2. (ii)

    Find the probability that the "Treasurer" and "Secretary" are both girls.

Work through it with AI →

It won't just hand you the answer — it asks where you're stuck.How the checker works

Worked solution

7(i)

1. Calculate the probability that all 4 selected committee members are girls.

P(all girls)=1124×1023×922×821=5161\mathrm{P}(\text{all girls}) = \frac{11}{24} \times \frac{10}{23} \times \frac{9}{22} \times \frac{8}{21} = \frac{5}{161}

2. Calculate the probability that all 4 selected committee members are boys.

P(all boys)=1324×1223×1122×1021=65966\mathrm{P}(\text{all boys}) = \frac{13}{24} \times \frac{12}{23} \times \frac{11}{22} \times \frac{10}{21} = \frac{65}{966}

3. Use the complement rule to find the probability of having at least one girl and at least one boy.

P(at least 1 girl and 1 boy)=1−P(all girls)−P(all boys)=1−5161−65966=1−30966−65966=1−95966=871966\mathrm{P}(\text{at least 1 girl and 1 boy}) = 1 - \mathrm{P}(\text{all girls}) - \mathrm{P}(\text{all boys}) \\ = 1 - \frac{5}{161} - \frac{65}{966} \\ = 1 - \frac{30}{966} - \frac{65}{966} \\ = 1 - \frac{95}{966} = \frac{871}{966}

Answer

871966\frac{871}{966}

7(ii)

1. Set up the probability expression for the two specific roles being filled by girls

Since the remaining two roles (Chairperson and Vice Chairperson) can be filled by anyone and do not restrict the Treasurer and Secretary positions, we can focus entirely on picking 2 girls from the class for these 2 specific roles.

P(Treasurer and Secretary are girls)=P(1st is a girl)×P(2nd is a girl given 1st is a girl)\text{P(Treasurer and Secretary are girls)} = \text{P(1st is a girl)} \times \text{P(2nd is a girl given 1st is a girl)}
=1124×1023= \frac{11}{24} \times \frac{10}{23}

2. Evaluate and simplify the fraction

=110552=55276= \frac{110}{552} \\ = \frac{55}{276}
Answer

55276\frac{55}{276}

AI analysis — may contain mistakes.

Worked this one out on paper? Snap your steps and Gigle's AI math checker will circle the step where it went wrong.

More questions on this topic

  1. 1An amateur music composer is arranging a sequence of four musical notes followed by three beats. There are 7 possible no…
  2. 2Two brothers Kai and Leo are duathlon (running and cycling) athletes who train regularly. For each day, the probability…
  3. 3Colour Vision Deficiency is the decreased ability to see differences in colours. Someone with normal colour vision will…
  4. 4A pair of siblings, Abbey and Betty, went to a book fair with six other friends. The eight of them stood in a queue in r…
  5. 5A particular disease is known to infect approximately 1 in every 10 000 people. A test for this disease is 99.9% accurat…